# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 5.4 Bookkeeping: names versus interpretations
To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes, e.g. $\mathsf{GBC}+\mathsf{ETR}$ (optionally with Global Choice). Its set universe is denoted $V$ and satisfies $\mathsf{ZFC}$. All symmetric models $\mathcal{N}$, intermediate stages $M_{\alpha}$, and the final model $\mathcal{M}$ are constructed inside this set universe; only the bookkeeping map $\mathcal{B}$ and the iteration itself are treated as class objects in the ambient background.
@@
In a class-length iteration, the interpretation of a fixed name can change as the iteration proceeds. Therefore, we bookkeep names (codes) rather than interpreted objects: at stage $\alpha$ we take the $\alpha$-th code $\dot{f}$ from our fixed enumeration, and in the current intermediate model $M_{\alpha}$ we check whether $\dot{f}$ (as currently interpreted) is a surjection instance of the required type. If it is not, we force trivially at that stage (or include only the other package factors scheduled for $\alpha$). This is the standard bookkeeping convention for class-length forcing/iterations.
@@
At stage $\alpha$, let $\mathcal{B}(\alpha)=(\beta,c)$. We test whether the code $c$:
@@
    decodes to a valid $\mathbb{P}_{\alpha}$-name $\dot{f}$,
@@
    is hereditarily symmetric for the stage-$\alpha$ symmetric system $(\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha})$, and
@@
    interprets in $M_{\alpha}$ to a surjection of one of the following types:
@@
      $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$ (for $\mathsf{AC}_{\mathsf{WO}}$-packages).
If all conditions are met and the surjection $f$ does not already split in $M_{\alpha}$ (i.e., no section/right-inverse for $f$ exists yet), the successor forcing at stage $\alpha+1$ includes the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ (Definition 5.5); otherwise we force trivially at that code.
This arrangement ensures that every code is tested at unboundedly many stages, so a name that becomes a valid HS surjection at some stage $\beta$ will eventually be scheduled at some stage $\alpha>\beta$.
@@
For the $\mathsf{AC}_{\mathsf{WO}}$-component we schedule splittings for surjections $f:Y\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)$ (cf. Remark 4.22). This is the natural bookkeeping horizon: $\aleph^{*}(S)$ is the first ordinal $\theta$ (in the ambient stage model under discussion) such that there is no surjection $S\twoheadrightarrow\theta$, so $\lambda<\aleph^{*}(S)$ is exactly the range of well-ordered indices reachable from $S$-parameterized data.
@@
    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.30), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
@@ 5.5 The symmetric iteration (black box)
We now invoke the iteration framework via Remark 2.1. For the global conclusion (a final model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$), we take a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$. Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
At a successor stage $\alpha+1$, we consult the finite list $\mathcal{B}(\alpha)$ from Definition 5.59. In the current model $M_{\alpha}$, each code on this list either interprets to a genuine package instance (yielding a corresponding orbit package forcing $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ as in Definition 5.5), or it does not (in which case we take the trivial forcing for that code). Let $\dot{\mathbb{S}}_{\alpha}$ be a $\mathbb{P}_{\alpha}$-name for the finite-support product of the resulting orbit package posets. Then $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$.
At a successor stage $\alpha+1$, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the filter definition in Definition 5.48. Concretely, $\mathscr{G}_{\alpha+1}$ is generated by standard lifts of $\mathscr{G}_{\alpha}$ together with the (stage-$\alpha$) package automorphisms and the successor diagonal kernel elements $\kappa_{\alpha+1}(\pi,D^{\mathrm{new}})$ (Definition 5.23) acting on the new package coordinates, and $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pushforwards from $\tilde{\mathscr{F}}^{*}_{\alpha}$ together with the admissible subgroups $\Delta^{\uparrow}_{\alpha+1}(E,D)$. In particular, for each active package coordinate $d$ at stage $\alpha+1$ with countable support witnessed by $E_{d}$, the generator $\Delta^{\uparrow}_{\alpha+1}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.45 and 5.46.
Thus each successor filter explicitly includes the diagonal subgroups protecting the representative coordinates of the active packages.
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [5, Thm. 5.5]). We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@
Unlike ordinary forcing iterations in $\mathsf{ZFC}$, there is no general theorem showing that an iteration of symmetric extensions can always be presented as a single symmetric extension of the ground model. Indeed, the usual “two-step collapse” phenomena from forcing do not directly apply in the symmetric context, and the literature discusses serious obstacles to such a reduction. It remains open in general to what extent arbitrary symmetric iterations can be compressed to a single symmetric extension; see the discussion in [9, 18].
@@ 5.6 What the iteration yields
Work in the final symmetric model $\mathcal{M}$. Let $f:Y\twoheadrightarrow X$ be a surjection in $\mathcal{M}$ with $X,Y\subseteq T$.
Choose a hereditarily symmetric name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$. Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{f}$ (API item (e)), and let $\alpha:=\sup(I)+1$. Then $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name and $f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}$. Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name, its interpretation is unchanged by further forcing: $\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}$, so $f=f_{\alpha}$ in $\mathcal{M}$. Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$–name of bounded rank, and no forcing past stage $\alpha$ adds sets of that rank, Theorem 9.2 of Karagila [9] implies that $\dot{f}^{G_{\lambda}}=\dot{f}^{G_{\alpha}}$. Hence $f=f_{\alpha}$ in $\mathcal{M}$.
If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done. Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping.
By the bookkeeping (Definition 5.59), at some successor stage we schedule the orbit package $\mathbb{Q}_{[f_{\alpha}]}$. Let $\dot{s}_{f_{\alpha}}$ be the canonical $\mathbb{Q}_{[f_{\alpha}]}$-name for the generic section in the $f_{\alpha}$-coordinate. By Lemma 5.45, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ in $M_{\beta}$. By Lemma 5.47, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence $s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$.
@@
At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support that is cofinal in $\lambda$, so stage-bounding arguments do not apply. The iteration machinery of [5] handles such limits using limit-filter $\omega_{1}$-completeness together with the appropriate block/factorization analysis (cf. [5, Rem. 3.9, Thm. 5.5]), and we rely on those limit-stage conclusions.
@@
Since $E$ is countable and $\omega_{1}$ is uncountable, there exist distinct $\beta,\gamma\in\omega_{1}\setminus E$. Let $\pi=(\beta\ \gamma)\in\operatorname{Sym}(\omega_{1})$ be the transposition swapping $\beta$ and $\gamma$. Then $\pi\in\operatorname{Fix}(E)$.
@@
We argue by induction along the symmetric iteration. The base model $M_{0}=\mathcal{N}$ satisfies $\mathsf{DC}$ by Theorem 3.11.
For the successor step, assume $M_{\alpha}\models\mathsf{ZF}+\mathsf{DC}$. The successor stage is obtained by forcing with the stage symmetric system at $\alpha$ and then taking the hereditarily symmetric submodel $M_{\alpha+1}$.
By Lemma 5.6, each orbit package $\mathbb{S}_{[f]}$ used at stage $\alpha$ is $\omega_{1}$-closed (and hence adds no new $\omega$-sequences of ordinals), and the finite-support product $\dot{\mathbb{S}}_{\alpha}$ is again $\omega_{1}$-closed for the same reason. By Lemma 5.16, the relevant stage filter is $\omega_{1}$-complete. Therefore the $\mathsf{DC}$-preservation theorem for symmetric extensions applies (API item (c), cf. [5, Thm. 3.20]), yielding $M_{\alpha+1}\models\mathsf{DC}$.
At limit stages, $\mathsf{DC}$ is preserved by the countable-support symmetric iteration machinery (API item (c), and the limit-filter generation/completeness from API item (d); cf. [5, Thm. 3.20] and [5, Thm. 5.5] for the limit construction).
Thus every stage model satisfies $\mathsf{DC}$, in particular the final model $\mathcal{M}\models\mathsf{DC}$. ∎
Definition 5.66 (Supported subforcing and supported names).
Fix a stage $\lambda$ and a globally admissible pair $(E,D)$ at stage $\lambda$. Define
$\mathbb{P}_{\lambda}\restriction(E,D)=\{p\in\mathbb{P}_{\lambda}:p\restriction 0\in\operatorname{Add}(\omega,E)\ \text{and}\ \operatorname{supp}_{>0}(p)\subseteq D\},$
where $\operatorname{supp}_{>0}(p):=\{c\in\mathrm{Coords}_{<\lambda}:p(c)\neq 1_{c}\}$ is the set of nontrivial higher-stage coordinates of $p$.
A $\mathbb{P}_{\lambda}$-name $\dot{x}$ is $(E,D)$-supported if every condition appearing in the transitive closure $\operatorname{tc}(\dot{x})$ belongs to $\mathbb{P}_{\lambda}\restriction(E,D)$.
Lemma 5.67 (Generic restriction to a supported subforcing).
Let $(E,D)$ be globally admissible at stage $\lambda$, and let $\mathbb{Q}=\mathbb{P}_{\lambda}\restriction(E,D)$. Then every maximal antichain $A\subseteq\mathbb{Q}$ is also a maximal antichain in $\mathbb{P}_{\lambda}$. Consequently, if $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ is $V$-generic, then $G_{E,D}:=G_{\lambda}\cap\mathbb{Q}$ is $V$-generic for $\mathbb{Q}$, and $V[G_{E,D}]\subseteq V[G_{\lambda}]$.
Proof.
Let $A\subseteq\mathbb{Q}$ be maximal, and let $p\in\mathbb{P}_{\lambda}$. Put $p^{\prime}:=p\restriction(E,D)\in\mathbb{Q}$. Choose $q\in A$ compatible with $p^{\prime}$ in $\mathbb{Q}$, and let $r\in\mathbb{Q}$ witness this, i.e. $r\leq_{\mathbb{Q}}q,p^{\prime}$. Since $p$ extends $p^{\prime}$ on the coordinates that $r$ constrains, and $r$ is trivial off $E\cup D$, we can amalgamate $r$ with $p$ to obtain a common extension in $\mathbb{P}_{\lambda}$. Thus $q$ is compatible with $p$ in $\mathbb{P}_{\lambda}$, so $A$ is predense in $\mathbb{P}_{\lambda}$ and hence maximal there. Now $G_{\lambda}$ meets every maximal antichain in $\mathbb{P}_{\lambda}$, so $G_{\lambda}\cap A\neq\emptyset$. Since $A\subseteq\mathbb{Q}$, this gives $(G_{\lambda}\cap\mathbb{Q})\cap A\neq\emptyset$, proving $\mathbb{Q}$-genericity. ∎
Lemma 5.68 (Supported names are $\mathbb{Q}$-names and evaluate correctly).
Let $(E,D)$ be globally admissible at stage $\lambda$, let $\mathbb{Q}=\mathbb{P}_{\lambda}\restriction(E,D)$, and let $G_{\lambda}\subseteq\mathbb{P}_{\lambda}$ be $V$-generic with $G_{E,D}=G_{\lambda}\cap\mathbb{Q}$. If $\dot{x}$ is $(E,D)$-supported, then $\dot{x}$ is (canonically) a $\mathbb{Q}$-name and
$\dot{x}^{G_{\lambda}}=\dot{x}^{G_{E,D}}.$
Proof.
Since every condition appearing in $\operatorname{tc}(\dot{x})$ lies in $\mathbb{Q}$, we may view $\dot{x}$ as a $\mathbb{Q}$-name (by the usual identification of names with the same transitive closure). The valuation equality is proved by induction on name-rank: for each pair $(\dot{y},p)\in\dot{x}$ we have $p\in\mathbb{Q}$, hence $p\in G_{\lambda}$ iff $p\in G_{E,D}$, and the induction hypothesis applies to $\dot{y}$. ∎
Lemma 5.69 (Triviality on supported names).
If $\dot{x}$ is $(E,D)$-supported and $g\in\Delta_{\lambda}^{\uparrow}(E,D)$, then $g\cdot\dot{x}=\dot{x}$.
Proof.
By induction on $\operatorname{rank}(\dot{x})$. If $p\in\mathbb{P}_{\lambda}\restriction(E,D)$, then $g(p)=p$ since the stage–0 projection of $g$ fixes $E$ pointwise and $g$ fixes every coordinate in $D$ pointwise. Now apply the recursive definition of the action on names. ∎

### v4

@@ 5.4 Bookkeeping: names versus interpretations
To speak formally about an $\mathrm{Ord}$-length bookkeeping function and an $\mathrm{Ord}$-length symmetric iteration as class objects, we work in a background theory of classes supporting definable class recursion for the iteration, e.g. $\mathsf{GBC}+\mathsf{ETR}$, and we fix a class well-order of $V$ (equivalently, assume Global Choice in the background). Its set universe is denoted $V$ and satisfies $\mathsf{ZFC}$. All symmetric models $\mathcal{N}$, intermediate stages $M_{\alpha}$, and the final model $\mathcal{M}$ are constructed inside this set universe; only the bookkeeping map $\mathcal{B}$ and the iteration itself are treated as class objects in the ambient background.
@@
In a class-length iteration, the interpretation of a fixed name can change as the iteration proceeds. Therefore, we bookkeep names (codes) rather than interpreted objects: at stage $\alpha$ we take the $\alpha$-th code $\dot{f}$ from our fixed enumeration, and in the current intermediate model $M_{\alpha}$ we check whether $\dot{f}$ (as currently interpreted) is a surjection instance of the required type. If it is not, we force trivially at that stage (or include only the other package factors scheduled for $\alpha$). We also use the canonical inclusion of $\mathbb{P}_{\beta}$-names into $\mathbb{P}_{\alpha}$-names for $\beta<\alpha$, so a code intended for an earlier stage can be retested later. This is the standard bookkeeping convention for class-length forcing/iterations.
@@
At stage $\alpha$, let $\mathcal{B}(\alpha)=(\beta,c)$. We attempt to decode $c$ as a $\mathbb{P}_{\beta}$-name $\dot{f}$ (in some fixed coding of names), and when $\beta\leq\alpha$ we view $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name via the canonical inclusion of $\mathbb{P}_{\beta}$-names into $\mathbb{P}_{\alpha}$-names. We then test whether (in the current stage model $M_{\alpha}$):
@@
    $\dot{f}$ is defined (the decoding succeeds) and $\beta\leq\alpha$,
@@
    $\dot{f}\in\operatorname{HS}_{\alpha}$ for the stage-$\alpha$ symmetric system $(\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha})$, and
@@
    $f:=\dot{f}^{G_{\alpha}}$ is a surjection instance of one of the following types:
@@
      $f:S\times\eta\twoheadrightarrow\lambda$ with $\eta,\lambda\in\mathrm{Ord}$ (for $\mathsf{AC}_{\mathsf{WO}}$-packages),
and moreover $f$ does not already split in $M_{\alpha}$ (i.e., no section/right-inverse for $f$ exists in $M_{\alpha}$).
If the test succeeds, we define the stage-$\alpha$ iterand $\dot{\mathbb{Q}}_{\alpha}$ so that in $M_{\alpha}$ it is the corresponding orbit package $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ (Definition 5.5); otherwise we set $\dot{\mathbb{Q}}_{\alpha}$ to be the trivial forcing. Thus $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}$.
This arrangement ensures that every code is tested at unboundedly many stages, so a name that becomes a valid HS surjection instance at some stage $\beta$ will eventually be scheduled at some stage $\alpha>\beta$.
@@
For the $\mathsf{AC}_{\mathsf{WO}}$-component we schedule splittings for surjections $f:S\times\eta\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)$ (cf. Remark 4.20). This is the natural bookkeeping horizon: $\aleph^{*}(S)$ is the first ordinal $\theta$ (in the ambient stage model under discussion) such that there is no surjection $S\twoheadrightarrow\theta$, so $\lambda<\aleph^{*}(S)$ is exactly the range of well-ordered indices reachable from $S$-parameterized data.
@@
    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.17 and 4.28), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
@@ 5.5 The symmetric iteration (black box)
We now invoke the iteration framework via Remark 2.4. Working in the background metatheory, we form a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$. Concretely, we build an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §3.6], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$. The next subsection states the properties of the final symmetric model that we prove.
At a successor stage $\alpha+1$, we consult the bookkeeping output $\mathcal{B}(\alpha)=(\beta,c)$ from Definition 5.61. In the current model $M_{\alpha}$, the code $c$ either decodes (via the canonical inclusion $\mathbb{P}_{\beta}$-names $\subseteq$ $\mathbb{P}_{\alpha}$-names) to a genuine package instance, yielding a corresponding orbit package forcing $\mathbb{Q}_{[f]}$ or $\mathbb{R}_{[f]}$ as in Definition 5.5, or it does not, in which case we take trivial forcing for that stage. Let $\dot{\mathbb{S}}_{\alpha}$ be the resulting $\mathbb{P}_{\alpha}$-name for the stage iterand (orbit package or trivial). Then $\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{S}}_{\alpha}$.
At successor stages, the group $\mathscr{G}_{\alpha+1}$ and filter $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ are those determined by the symmetric-iteration framework together with the diagonal infrastructure from Subsection 5.2 and the modified stage filters from Definition 5.50. Concretely, $\tilde{\mathscr{F}}^{*}_{\alpha+1}$ is the normal filter generated by pullbacks $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$, together with admissible subgroups $\Delta_{\alpha+1}^{\uparrow}(E,D)$. In particular, for each active package coordinate $d$ at stage $\alpha+1$ whose countable support is witnessed by $E_{d}$, the generator $\Delta_{\alpha+1}^{\uparrow}(E_{d},\{d\})$ belongs to $\tilde{\mathscr{F}}^{*}_{\alpha+1}$, so the canonical package generics are hereditarily symmetric by Lemmas 5.47 and 5.48.
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group as provided by the symmetric-iteration framework (Remark 2.4(a),(b)). The limit-stage filter is $\tilde{\mathscr{F}}^{*}_{\lambda}$ as defined in Definition 5.50, and its normality and $\omega_{1}$-completeness are recorded in Lemma 5.54. For set-localization in the $\mathrm{Ord}$-length model we use only the stage-support lemma (Lemma 5.64) below. We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@
Unlike ordinary forcing iterations in $\mathsf{ZFC}$, there is no general theorem showing that an iteration of symmetric extensions can always be presented as a single symmetric extension of the ground model. Indeed, the usual “two-step collapse” phenomena from forcing do not directly apply in the symmetric context, and the literature discusses serious obstacles to such a reduction. It remains open in general to what extent arbitrary symmetric iterations can be compressed to a single symmetric extension; see the discussion in [9].
@@ 5.6 What the iteration yields
Lemma 5.64 (Stage-localization of set-names at class length).
Let $\dot{x}$ be a set-sized $\mathbb{P}_{\mathrm{Ord}}$-name, and define its stage-support
$I(\dot{x})\ :=\ \bigcup\{\operatorname{supp}(p)\cup I(\dot{y}):\langle\dot{y},p\rangle\in\dot{x}\},$
where $\operatorname{supp}(p)\subseteq\mathrm{Ord}$ is the (countable) support of the condition $p$. Then $I(\dot{x})$ is a set of ordinals, hence $\alpha:=\sup(I(\dot{x}))+1<\mathrm{Ord}$. Moreover:
  1. 1.
    $\dot{x}$ is in fact a $\mathbb{P}_{\alpha}$-name (all conditions appearing in $\dot{x}$ lie in $\mathbb{P}_{\alpha}$),
  2. 2.
    for any $V$-generic $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ and $G_{\alpha}:=G_{\infty}\cap\mathbb{P}_{\alpha}$, we have $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$.
Proof.
Define $I(\dot{x})$ by recursion on the rank of the name $\dot{x}$. Since $\dot{x}$ is set-sized, Replacement and Union yield that $I(\dot{x})$ is a set of ordinals, hence bounded in $\mathrm{Ord}$. Let $\alpha=\sup(I(\dot{x}))+1$. Every condition $p$ appearing in $\dot{x}$ has $\operatorname{supp}(p)\subseteq I(\dot{x})\subseteq\alpha$, so $p\in\mathbb{P}_{\alpha}$, hence $\dot{x}$ is a $\mathbb{P}_{\alpha}$-name. The valuation recursion for $\dot{x}^{G_{\infty}}$ consults only conditions appearing in $\dot{x}$, all of which lie in $\mathbb{P}_{\alpha}$, hence depends only on $G_{\alpha}$. Therefore $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\alpha}}$. ∎
@@
Work in the final symmetric model $\mathcal{M}$. Recall Definition 4.8. Let $f:Y\twoheadrightarrow X$ be a surjection in $\mathcal{M}$ with $X,Y\subseteq T$.
Choose a hereditarily symmetric set-name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$. By Lemma 5.64, letting $\alpha:=\sup(I(\dot{f}))+1$, we may view $\dot{f}$ as a $\mathbb{P}_{\alpha}$-name and we have
$f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}\qquad\text{and}\qquad\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}.$
Thus $f=f_{\alpha}\in M_{\alpha}\subseteq\mathcal{M}$.
If $f_{\alpha}$ already has a section in $M_{\alpha}$, then $f$ splits in $\mathcal{M}$ and we are done. Otherwise, $f_{\alpha}$ is eligible for scheduling by the bookkeeping. By the final paragraph of Definition 5.61, there is some stage $\beta\geq\alpha$ such that either $f_{\alpha}$ already splits in $M_{\beta}$ (in which case we are done), or else the successor iterand at stage $\beta$ is the orbit $\mathsf{PP}$-package $\mathbb{Q}_{[f_{\alpha}]}$.
Assume we are in the latter case, and fix such a stage $\beta$. Let $\dot{s}_{f_{\alpha}}$ be the canonical $\mathbb{Q}_{[f_{\alpha}]}$-name for the generic section in the $f_{\alpha}$-coordinate. In the full package extension, this coordinate generic yields a section of $f_{\alpha}$ (cf. Proposition 4.15).
By Lemma 5.47, $\dot{s}_{f_{\alpha}}$ is hereditarily symmetric for the package symmetric system $(\mathbb{Q}_{[f_{\alpha}]},\mathscr{G}_{[f_{\alpha}]},\mathscr{F}_{[f_{\alpha}]})$ over $M_{\beta}$. By Lemma 5.49, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{s}_{f_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence
$s_{f_{\alpha}}:=\dot{s}_{f_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}.$
@@
At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support cofinal in $\lambda$, so the usual stage-bounding arguments do not apply. In this manuscript we handle such limits by using the modified limit filter (Definition 5.50), whose normality and $\omega_{1}$-completeness are recorded in Lemma 5.54, together with the core admissibility lemma (Lemma 5.52). This matches the treatment of countable-cofinality limits in the general iteration framework ([5, §3.5–§3.6]).
@@
Since $E$ is countable, $\delta:=\sup(E)<\omega_{1}$. Choose distinct $\beta,\gamma\in\omega_{1}$ with $\delta<\beta<\gamma$. Then $\beta,\gamma\notin E$. Let $\pi=(\beta\ \gamma)\in\operatorname{Sym}(\omega_{1})$ be the transposition swapping $\beta$ and $\gamma$. Then $\pi\in\operatorname{Fix}(E)$.
@@
Let $A$ be a set in $\mathcal{M}$ and let $R\subseteq A\times A$ be a relation in $\mathcal{M}$ such that $\forall x\in A\,\exists y\in A\ (xRy)$. We show there is an $\omega$-sequence $\langle a_{n}:n<\omega\rangle$ in $\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$.
Choose set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-names $\dot{A},\dot{R}\in\operatorname{HS}_{\mathrm{Ord}}$ such that $A=\dot{A}^{G_{\infty}}$ and $R=\dot{R}^{G_{\infty}}$. By Lemma 5.64, letting $\alpha:=\sup(I(\dot{A})\cup I(\dot{R}))+1$, we have $A,R\in M_{\alpha}$.
By Lemma 5.54(iv), the stage filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ is $\omega_{1}$-complete. Therefore, by Remark 2.4(d), $M_{\alpha}\models\mathsf{DC}$. Applying $\mathsf{DC}$ inside $M_{\alpha}$ to $(A,R)$, we obtain a sequence $\langle a_{n}:n<\omega\rangle\in M_{\alpha}\subseteq\mathcal{M}$ with $a_{n}Ra_{n+1}$ for all $n$. Hence $\mathcal{M}\models\mathsf{DC}$. ∎